# Enhancing Multimodal Large Language Models with Vision Detection Models: An Empirical Study

### v2

@@ 3.4 Quantitative Evaluation Settings
We employ 10 benchmarks to evaluate different MLLM capabilities in the main page: VQAv2 [14], GQA [18], and MME [12] measure comprehensive VQA capabilities; MMBench [33] and SEED-Bench [25] evaluate perceptual and reasoning abilities; TextVQA [46] assesses text recognition abilities; MM-Vet [56] evaluates abilities for managing complex task with local information; and POPE [28] measures fine-grained object localization abilities. Regarding the GQA benchmark, which tends to under-estimate performance due to its textual exact match evaluation setting, we utilize a subset that retains unambiguous questions, denoted as GQA*. Detailed information of the GQA* is provided in Appendix E.1.
For a more comprehensive and convenient comparison, we employ an aggregated score across all considered benchmarks. Specifically, we first normalize the score for each benchmark as $s_{norm}=(s-s_{min})/(s_{max}-s_{min})$, where $s$ denotes the original benchmark score, $s_{min}$ and $s_{max}$ refers to the lowest and highest score across all considered methods respectively. Subsequently, we compute the mean of $s_{norm}$ across all benchmarks and denote it as $\overline{s}_{norm}$.
@@ 4 Main Results and Analysis
4.1 Lesson 1: Original MLLMs Struggle with Comprehending Detection Information
Initially, we directly feed the detection information into the original LLaVA-1.5, aiming to observe whether the original LLaVA-1.5 can comprehend the detection information we input. We call this fusion strategy “Training-free Infusion” (TFI) as introduced in Section 3.3.
Table 1: Comparison between “Training-free Infusion” (TFI) models and LLaVA-1.5.
MLLM VQAv2 GQA* VQAT POPE MMEP MMEC MMB MMBCN MM-Vet SEED LLaVA-1.5-7B 78.5 79.6 58.2 85.9 1510.7 355.7 64.3 58.3 30.5 58.6 TFI-7B 78.5 79.2 $\downarrow$ 59.2 $\uparrow$ 89.9 $\uparrow$ 1497.0 $\downarrow$ 401.0 $\uparrow$ 65.0 $\uparrow$ 57.2 $\downarrow$ 33.7 $\uparrow$ 60.6 $\uparrow$ LLaVA-1.5-13B 80.0 81.0 61.3 85.9 1531.3 295.4 67.7 63.6 35.4 61.6 TFI-13B 76.6 $\downarrow$ 79.0 $\downarrow$ 59.6 $\downarrow$ 88.3 $\uparrow$ 1453.6 $\downarrow$ 401.0 $\uparrow$ 65.0 $\downarrow$ 57.5 $\downarrow$ 31.7 $\downarrow$ 60.7 $\downarrow$
@@ 4.1 Lesson 1: Original MLLMs Struggle with Comprehending Detection Information
The results are presented in Table 1. We can see that TFI-7B exhibits partial enhancement in some benchmarks, while TFI-13B shows a discernible decline. Both models show significant improvement on the POPE benchmark, which evaluates object hallucination, indicating that the infused object detection information is helpful for object detection tasks. Additionally, they exhibit robust performance on the MME-Cognition benchmark, which contains numerous questions related to text within images, suggesting that the OCR information is also demonstrating efficacy.
@@
However, other benchmarks exhibit fluctuations in scores, implying a deficiency in our model’s utilization of detection information. Upon closer analysis, we believe that the infusion of detection information introduces extraneous content, which becomes noise, thereby adversely affecting the inference accuracy. Therefore, we proceed to train the model to help it better extract useful information from the detection information and eliminate noise.
4.2 Lesson 2: Detection Re-Training has Adverse Effects on Comprehending ViT Features
Researchers commonly employ in-context learning (ICL) [9] to instruct LLMs in the recognition of specific input formats. However, for MLLMs, ICL becomes difficult as it requires MLLMs to process lengthy input sequences combining image features in tokens format. Therefore, we opt against using ICL for MLLMs in favor of a training-based approach, and shifting into the second fusion strategy introduced in Section 3.3, “Retraining Based Infusion” (RBI). In this strategy, we retrain LLaVA-1.5 with its original training pipeline, concurrently infusing the detection information.
Table 2: Comparison between “Retraining Based Infusion” (RBI) models and LLaVA-1.5.
MLLM VQAv2 GQA* VQAT POPE MMEP MMEC MMB MMBCN MM-Vet SEED
LLaVA-1.5-7B 78.5 79.6 58.2 85.9 1510.7 355.7 64.3 58.3 30.5 58.6
RBI-7B 78.5 76.6 $\downarrow$ 60.0 $\uparrow$ 89.3 $\uparrow$ 1454.5 $\downarrow$ 412.0 $\uparrow$ 66.2 $\uparrow$ 60.6 $\uparrow$ 31.5 $\uparrow$ 60.8 $\uparrow$
LLaVA-1.5-13B 80.0 81.0 61.3 85.9 1531.3 295.4 67.7 63.6 35.4 61.6
RBI-13B 79.2 $\downarrow$ 78.0 $\downarrow$ 61.7 $\uparrow$ 89.2 $\uparrow$ 1491.2 $\downarrow$ 409.6 $\uparrow$ 69.5 $\uparrow$ 63.2 $\downarrow$ 35.1 $\downarrow$ 62.5 $\uparrow$
@@ 4.2 Lesson 2: Detection Re-Training has Adverse Effects on Comprehending ViT Features
RBI-13B w/o DI 77.3 $\downarrow$ 76.0 $\downarrow$ 58.2 $\downarrow$ 83.4 $\downarrow$ 1442.6 $\downarrow$ 310.7 $\uparrow$ 68.5 $\uparrow$ 61.7 $\downarrow$ 30.6 $\downarrow$ 61.6 $\uparrow$
Performance Improvement Relative to LLaVA-1.5 and TFI.
[…]
Adverse Impact of RBI on ViT Feature Comprehension.
RBI models do not show improvement across all benchmarks. On VQAv2, GQA*, and MME, there is a noticeable decline. We speculate that the unexpected results are due to an over-infusion that negatively affects the model’s ability to learn how to utilize features from ViT. Thus, we then evaluate the performance of RBI models with no detection information applied. In this way, their benchmark scores are only related to ViT features. As shown in Table 3, the current models show a noticeable performance lag compared to LLaVA-1.5, suggesting that RBI strategy does impact the MLLM in learning how to use the image features extracted from ViT, calling for further improved fusion strategies.
Table 4: Comparison between “Retraining Based Infusion” (RBI) models, “Fine-tuning Based Infusion” (FTBI) models, and SOTA methods on 10 benchmarks. Our FTBI approach effectively infuses detection information into LLaVA-1.5, leading to a considerable performance enhancement across all 10 benchmarks. $\overline{s}_{norm}$ is an aggregated performance measure across all considered benchmarks. Bold and underlined results indicate the best and second-best performance respectively. MME represents the summation of MMEP and MMEC, with detailed information in Appendix E.2.
MLLM VQAv2 GQA* VQAT POPE MME MMB MMBCN MM-Vet SEED $\overline{s}_{norm}$
BLIP-2 (13B) 41.0 41.0 42.5 85.3 1583.8 - - 22.4 46.4 -
InstructBLIP (13B) 75.3 64.3 50.7 78.9 1504.6 34.5 15.5 25.6 43.6 0.245
Qwen-VL (7B) 78.8 74.8 63.8 78.1 1667.0 38.2 7.4 38.2 56.3 0.512
Qwen-VL-Chat (7B) 78.2 77.9 61.5 79.2 1848.2 60.6 56.7 41.5 58.2 0.756
SPHINX (13B) 78.1 80.4 51.6 80.7 1798.3 66.9 56.2 36.0 69.1 0.746
LLaVA-1.5-7B 78.5 79.6 58.2 85.9 1866.4 64.3 58.3 30.5 58.6 0.767
RBI-7B (our) 78.5 76.6 $\downarrow$ 60.0 $\uparrow$ 89.3 $\uparrow$ 1866.5 $\uparrow$ 66.2 $\uparrow$ 60.6 $\uparrow$ 31.5 $\uparrow$ 60.8 $\uparrow$ 0.828 $\uparrow$
FTBI-7B (our) 79.0 $\uparrow$ 80.1 $\uparrow$ 60.1 $\uparrow$ 88.9 $\uparrow$ 1880.5 $\uparrow$ 67.3 $\uparrow$ 60.2 $\uparrow$ 35.2 $\uparrow$ 60.8 $\uparrow$ 0.863 $\uparrow$
LLaVA-1.5-13B 80.0 81.0 61.3 85.9 1826.7 67.7 63.6 35.4 61.6 0.843
RBI-13B (our) 79.2 $\downarrow$ 78.0 $\downarrow$ 61.7 $\uparrow$ 89.2 $\uparrow$ 1900.9 $\uparrow$ 69.5 $\uparrow$ 63.2 $\downarrow$ 35.1 $\downarrow$ 62.5 $\uparrow$ 0.894 $\uparrow$
FTBI-13B (our) 80.3 $\uparrow$ 81.8 $\uparrow$ 61.8 $\uparrow$ 88.8 $\uparrow$ 1920.5 $\uparrow$ 71.4 $\uparrow$ 65.2 $\uparrow$ 38.9 $\uparrow$ 62.3 $\uparrow$ 0.940 $\uparrow$
@@
4.3 Lesson 3: Suitable Fine-tuning Achieves Good Trade-off for Detection Infusion
To enhance the model’s ability to make well-informed trade-offs between ViT features and detection information, for the third fusion strategy, we leverage the well-trained parameters of LLaVA-1.5 and fine-tune our models based on it. We call this fusion strategy “Fine-tuning Based Infusion”, abbreviated as FTBI.
Performance Improvement Relative to LLaVA-1.5 and RBI.
@@ 4.3 Lesson 3: Suitable Fine-tuning Achieves Good Trade-off for Detection Infusion
As shown in Table 4, both the 7B and 13B models of FTBI exhibit superior performance compared to LLaVA-1.5. Simultaneously, as indicated in Appendix C.2, when detection information is not infused, FTBI models demonstrate significant improvement in comparison to the RBI models, indicating that the fine-tuning strategy effectively makes trade-offs between ViT features and detection information.
@@
Upon detailed analysis, we find that FTBI models exhibit a visible improvement on comprehensive VQA benchmarks such as VQAv2, GQA*, and MME. On the benchmarks that evaluate perceptual and reasoning abilities, such as MMBench and SEED-Bench, the models’ performance undergoes a noticeable enhancement. Moreover, the infusion of object detection information significantly improves performance on both the POPE benchmark, which evaluates object hallucinations, and the MM-Vet benchmark, which measures the ability to utilize local information. Due to the infusion of OCR information, the models also exhibit commendable performance on text-related benchmarks such as TextVQA and MME-cognition. Finally, on the aggregated performance measure $\overline{s}_{norm}$, FTBI models significantly outperform LLaVA-1.5 and other considered SOTA models, highlighting the substantial effectiveness of our approach.
Case Study.
In Figure 1, we list examples on which LLaVA-1.5 produces incorrect answers while our model yields correct responses (more examples are presented in Appendix A.1). With object detection, our model effectively achieves precise counts and delineates the respective positions of specified objects. With OCR, our model proficiently generates textual content in response to the localization requirements. Furthermore, our model can use the object detection information to refine its comprehension of the information conveyed by particular objects. For instance, when the model discerns the location of a clock within the image, it can more effectively extract the displayed content on the clock, thereby acquiring the correct time.
4.4 Lesson 4: Detection Models Can be Flexibly and Effectively Replaced
DINO is a closed-set object detection model and cannot deal with images containing uncommon objects or specific entities such as certain celebrities and objects with attributive modifiers. Fortunately, our model is modular, which means that we have the flexibility to replace the embedded detection models. Additionally, the modular structure allows our model to retain the training effectiveness of comprehending specially designed detection information after the module replacement.
Constructing Detection Information with Grounding DINO.
In this experiment, we will substitute the embedded closed-set detection model DINO with an open-set object detection model called Grounding DINO [32]. Grounding DINO is designed to detect objects related to user-input. With this model, our model can pinpoint the targets by referencing the object names mentioned in questions. To achieve this, we first extract target names from the input questions, and then combine them to create prompts. Grounding DINO then follows the prompts to generate location information for the targets. Finally, we convert the outputs into the specially designed detection information.
@@ 4.4 Lesson 4: Detection Models Can be Flexibly and Effectively Replaced
Figure 4: Examples on which LLaVA-1.5 fails while our model with open-set object detection information succeeds.
@@
In Table 3, we compare the performance of TFI-7B and FTBI-7B after replacing the detection model DINO with Grounding DINO. We use VQAv2, GQA*, POPE, MM-Vet, and SEED-Bench for evaluation as they contain questions from which effective noun phrases can be extracted. Due to the low detection accuracy of Grounding DINO, some noise is introduced, which results in lower evaluation scores for both models compared to LLaVA-1.5-7B. However, because FTBI-7B has been trained with DINO and it can filter out some noise, the performance of FTBI-7B is superior to that of TFI-7B. These results validate that the training effectiveness remains after we replace the detection model.
Case Study.
In Figure 4, we enumerate cases where LLaVA-1.5 provides incorrect answers while our model with open-set object detection information produces correct responses. Although Grounding DINO may output error messages, when its outputs are correct, this information can effectively assist our model. With open-set object detection information, our model can carry out counting tasks and localization tasks more purposefully, as it can leverage the target names in questions to get specific positional information.
4.5 Overview of More Experiments

### v3

@@ 3.4 Quantitative Evaluation Settings
We employ 10 widely recognized benchmarks to evaluate different MLLM capabilities: VQAv2 (Goyal et al., 2017), GQA (Hudson & Manning, 2019), and MME (Fu et al., 2023) measure comprehensive VQA capabilities; MMBench (Liu et al., 2023d) and SEED-Bench (Li et al., 2023a) evaluate perceptual and reasoning abilities; TextVQA (Singh et al., 2019) assesses text recognition abilities; MM-Vet (Yu et al., 2023) evaluates abilities for managing complex task with fine-grained image details; and POPE (Li et al., 2023e) measures fine-grained object localization abilities. It’s noteworthy that we evaluate the models using a subset of GQA benchmark, denoted as GQA*, which retains unambiguous questions. Detailed information of the GQA* is provided in Appendix E.1. For a more comprehensive and convenient comparison, we compute the average percentage improvement of the models, trained with different strategies, over the original models across the 10 benchmarks, denoted as $\Delta$.
@@ 4 Main Results and Analysis
4.1 Overview and Organization
In this section, we systematically evaluate the performance improvements of the enhanced MLLMs over the original models under various training strategies. We find that the FTBI training strategy yields the best results. As shown in Table 1, on 10 well-recognized MLLM benchmarks, FTBI-7B and FTBI-13B exhibits a 3.99% and 3.30% improvement compared to LLaVA-1.5-7B and LLaVA-1.5-13B respectively. Besides, FTBI-13B outperforms TFI-13B by 6.71%.
We will delve into the progressive exploration of the studied training strategies (TFI in Section 4.2, RBI in Section 4.3, and FTBI in Section 4.4). Additionally, in Section 4.5, we will test the substitution of the deployed object detection model and explore whether the fine-tuned MLLM can retain its training effects after the replacement. Moreover, in Appendix D, we will provide further experimental analysis with a new MLLM, Qwen-VL, and a new detector, YOLOv8.
Table 1: Comparison between “Training-free Infusion”(TFI) models, “Retraining Based Infusion” (RBI) models, “Fine-tuning Based Infusion” (FTBI) models, and the original LLaVA-1.5 on 10 benchmarks. $\Delta$ represents the average percentage improvement relative to the original models. Bold and underlined results indicate the best and second-best performance respectively. MME represents the summation of MMEP and MMEC, with detailed information in Appendix E.2.
MLLM VQAv2 GQA* VQAT POPE MME MMB MMBCN MM-Vet SEED $\Delta$
LLaVA-1.5-7B 78.5 79.6 58.2 85.9 1866.4 64.3 58.3 30.5 58.6 -
TFI-7B 78.5 = 79.2 $\downarrow$ 59.2 $\uparrow$ 89.9 $\uparrow$ 1898.0 $\uparrow$ 65.0 $\uparrow$ 57.2 $\downarrow$ 33.7 $\uparrow$ 60.6 $\uparrow$ +2.30%
RBI-7B 78.5 = 76.6 $\downarrow$ 60.0 $\uparrow$ 89.3 $\uparrow$ 1866.5 $\uparrow$ 66.2 $\uparrow$ 60.6 $\uparrow$ 31.5 $\uparrow$ 60.8 $\uparrow$ +1.91%
FTBI-7B 79.0 $\uparrow$ 80.1 $\uparrow$ 60.1 $\uparrow$ 88.9 $\uparrow$ 1880.5 $\uparrow$ 67.3 $\uparrow$ 60.2 $\uparrow$ 35.2 $\uparrow$ 60.8 $\uparrow$ +3.99%
LLaVA-1.5-13B 80.0 81.0 61.3 85.9 1826.7 67.7 63.6 35.4 61.6 -
TFI-13B 76.6 $\downarrow$ 79.0 $\downarrow$ 59.6 $\downarrow$ 88.3 $\uparrow$ 1854.6 $\uparrow$ 65.0 $\downarrow$ 57.5 $\downarrow$ 31.7 $\downarrow$ 60.7 $\downarrow$ -3.41%
RBI-13B 79.2 $\downarrow$ 78.0 $\downarrow$ 61.7 $\uparrow$ 89.2 $\uparrow$ 1900.9 $\uparrow$ 69.5 $\uparrow$ 63.2 $\downarrow$ 35.1 $\downarrow$ 62.5 $\uparrow$ +0.72%
FTBI-13B 80.3 $\uparrow$ 81.8 $\uparrow$ 61.8 $\uparrow$ 88.8 $\uparrow$ 1920.5 $\uparrow$ 71.4 $\uparrow$ 65.2 $\uparrow$ 38.9 $\uparrow$ 62.3 $\uparrow$ +3.30%
4.2 Lesson 1: The Original MLLM Struggle with Comprehending Textual Detection Information
Initially, we input the textual detection information directly into the original LLaVA-1.5, aiming to observe whether it can comprehend and utilize this specially formatted information. We call this training strategy “Training-free Infusion” (TFI) as introduced in Section 3.3.
@@ 4.2 Lesson 1: The Original MLLM Struggle with Comprehending Textual Detection Information
The results are presented in Table 1. We can see that TFI-7B exhibits partial enhancement in some benchmarks, while TFI-13B shows a discernible decline. Both models show significant improvement on the POPE benchmark, which evaluates object hallucination, indicating that the infused object detection information works well. Besides, as shown in Appendix E.2, they exhibit robust performance on the MME-Cognition benchmark, which contains numerous questions related to text within images, suggesting that the OCR information is also demonstrating efficacy.
@@
However, other benchmark scores exhibit fluctuations, implying a deficiency in training-free models’ utilization of textual detection information. Upon closer analysis, we believe that the infusion of textual detection information introduces extraneous content, which may become noise, thereby adversely affecting the accuracy. In other words, if the models are not trained adaptively with the specially formatted detection information, it may not be able to effectively extract useful information from it and can be misguided by noise.
4.3 Lesson 2: Retraining has Adverse Effects on Comprehending ViT Features
In Section 4.2, we experimentally demonstrate that the studied MLLM with a training-free strategy fails to fully comprehend and use the textual detection information we input. Nevertheless, as demonstrated by numerous studies (Zhang et al., 2024b; Cabessa et al., 2024), adapting LLMs through training with specially formatted text helps them more effectively extract useful information from it, while identifying and filtering out noise within the text. Hence, we will then explore whether the retraining strategy can improve the model’s understanding of this textual detection information. For the “Retraining Based Infusion” (RBI) strategy, we retrain LLaVA-1.5 based on its original training pipeline, concurrently infusing the textual detection information.
Performance Improvement Relative to the Original Model.
[…]
@@ 4.3 Lesson 2: Retraining has Adverse Effects on Comprehending ViT Features
RBI-13B w/o DI 77.3 $\downarrow$ 76.0 $\downarrow$ 58.2 $\downarrow$ 83.4 $\downarrow$ 1442.6 $\downarrow$ 310.7 $\uparrow$ 68.5 $\uparrow$ 61.7 $\downarrow$ 30.6 $\downarrow$ 61.6
Adverse Impact of Retraining on ViT Feature Comprehension.
Nevertheless, RBI models do not show improvement across all benchmarks. While the 13B version of RBI shows a clear advantage over the training-free model, its improvement over the original model is still limited. Besides, the 7B version of RBI even performs similarly to the training-free model. These unexpected results may be due to the redundant information in the textual detection information, which negatively affects MLLM’s ability to learn how to utilize features from ViT (the image encoder) during training.
We then evaluate the performance of RBI models with no detection information applied during inference. Upon this, their benchmark scores are only related to ViT features. As shown in Table 2, the models show a noticeable performance lag compared to LLaVA-1.5, indicating that the retraining strategy does harm the model in learning how to use image features extracted from the image encoder. However, it is essential to note that the real world applications encompass a substantial amount of tasks that do not require fine-grained information but rather demand image-level information. Upon these tasks, the MLLM places greater reliance on ViT features. Therefore, while facilitating the model’s learning of how to utilize detection information, it is crucial to simultaneously ensure the model preserves its capability to leverage ViT features.
@@
4.4 Lesson 3: Suitable Fine-tuning Achieves Good Trade-offs between ViT Features and Textual Detection Information
As indicated in Section 4.3, retraining could inevitably pose challenges for MLLMs in precisely evaluating the significance of ViT features and detection information, leading to a decline in understanding ViT features and a decrease in performance on tasks unrelated to detection. For the third training strategy, we leverage the well-trained parameters of LLaVA-1.5. Specifically, we fine-tune the pre-trained LLaVA-1.5 for an additional epoch with the textual detection information infused, aiming to observe whether the fine-tuning strategy can enhance MLLMs’ ability to effectively balance between ViT features and detection information, and boost their performance on fine-grained image recognition. We call this training strategy “Fine-tuning Based Infusion”, abbreviated as FTBI.
Performance Improvement Relative to the Original Model, the Training-free Model, and the Retrained Model.
@@
As shown in Table 1, both the 7B and 13B versions of FTBI exhibit superior performance compared to LLaVA-1.5, TFI, and RBI, with the FTBI models outperform the original models by up to 3.99%, and surpass the training-free models by up to 6.71%. Simultaneously, as indicated in Table 3, when the detection information is not infused, FTBI models show significant improvement over the RBI models and achieve performance comparable to that of LLaVA-1.5, indicating that the fine-tuning strategy retains LLaVA-1.5’s original understanding of ViT features and effectively makes good trade-offs between ViT features and the detection information.
Table 3: If we do not infuse detection information to FTBI-7B and FTBI-13B during inference, their performance will be on par with LLaVA-1.5-7B and LLaVA-1.5-13B. “w/o DI” is an abbreviation for “without detection information.”
MLLM VQAv2 GQA* VQAT POPE MMEP MMEC MMB MMBCN MM-Vet SEED
LLaVA-1.5-7B 78.5 79.6 58.2 85.9 1510.7 355.7 64.3 58.3 30.5 58.6
RBI-7B w/o DI 76.4 74.8 56.6 85.5 1387.7 312.5 65.5 58.3 29.0 59.6
FTBI-7B w/o DI 78.0 78.4 57.1 86.0 1441.8 303.6 66.9 59.7 30.1 60.6
LLaVA-1.5-13B 80.0 81.0 61.3 85.9 1531.3 295.4 67.7 63.6 35.4 61.6
RBI-13B w/o DI 77.3 76.0 58.2 83.4 1442.6 310.7 68.5 61.7 30.6 61.6
FTBI-13B w/o DI 79.4 80.0 60.0 85.3 1525.7 320.0 70.8 64.8 36.0 61.7
@@
Upon detailed analysis on Table 1, we can find that FTBI models exhibit a visible improvement on comprehensive VQA benchmarks such as VQAv2, GQA*, and MME. On the benchmarks that evaluate perceptual and reasoning abilities, such as MMBench and SEED-Bench, the models’ performance undergoes a noticeable improvement. Moreover, the infusion of object detection information significantly improves performance on both the POPE benchmark, which evaluates object hallucinations, and the MM-Vet benchmark, which contains questions about fine-grained image recognition. Due to the infusion of OCR information, the models also exhibit commendable performance on text-related benchmarks such as TextVQA and MME-cognition. Finally, on the overall performance measure $\Delta$, FTBI models outperform LLaVA-1.5 by 3.99% and 3.30% for the 7B and 13B versions respectively. Besides, FTBI models outperform the TFI models by 1.69% and 6.71%, indicating that fine-tuning on textual detection information is effective and allows MLLMs to better comprehend and utilize the detection information.
Fine-tuned Models Can Still Perform Well Without Infusing Detection Information.
We assess the benchmark scores of FTBI models without infusing detection information during inference, aiming to evaluate their capacities in leveraging ViT features. The findings delineated in Table 3 demonstrate that the efficacy of FTBI models without detection information aligns closely with that of LLaVA-1.5, and they outperform RBI models without detection information across all benchmarks. It means that the fine-tuning strategy effectively empowers the model to assimilate and make use of image features extracted by ViT, suggesting that it achieves a good balance between image features and detection information. Therefore, the fine-tuning strategy is superior to the training-free strategy and the retraining strategy.
4.5 Lesson 4: Suitable Fine-Tuning Helps MLLMs Better Understand Specially Formatted Detection Information
In the previous experiments, we employ DINO to extract object detection information and successfully facilitate performance improvement for the MLLM. However, it is essential to note that DINO is a closed-set object detection model, capable of detecting only 80 common object categories. Nevertheless, images may contain uncommon objects or specific entities such as certain celebrities or objects with attributive modifiers. In such scenarios, the closed-set models exhibit limitations.
Fortunately, the studied MLLM architecture is modular, and the deployed detection models are independent of the MLLM. Hence, this architecture allows for flexible replacement of the deployed detection models. In this experiment, we will substitute the closed-set detector DINO with an open-set detector to observe whether, after the replacement, the finetuned MLLMs (FTBI) can still operate effectively and acquire broader capability for detection.
@@
Figure 4: Examples on which LLaVA-1.5 fails while the fine-tune model (FTBI-13B) with open-set object detection information succeeds.
Constructing Detection Information with Grounding DINO.
In this experiment, we substitute the embedded closed-set detector DINO with an open-set object detector called Grounding DINO (Liu et al., 2023c). Grounding DINO is designed to detect objects related to user-input. With this model, the studied MLLM can locate targets by referring to the object names mentioned in questions. To achieve this, we first extract target names from the input questions and combine them to create prompts. Grounding DINO then follows the prompts to generate location information for the targets. Finally, the outputs are converted into specially formatted detection information following the method in Section 3.1. Figure 3 shows an example of this process.
@@
In Table 3, we compare the performance of TFI-7B and FTBI-7B after replacing the detection model DINO with Grounding DINO. We use VQAv2, GQA*, POPE, MM-Vet, and SEED-Bench for evaluation as they contain questions from which effective object names can be extracted. Due to the low detection accuracy of Grounding DINO, some noise is introduced, which results in lower evaluation scores for both models compared to LLaVA-1.5-7B. However, as FTBI-7B has been fine-tuned with DINO and it can filter out some noise, the performance of FTBI-7B is superior to that of TFI-7B. These results validate that the training effect remains after we replace the detection model.
5 Overview of More Experiments
We list additional experimental details and parameter settings in the appendix and conduct further experiments to validate the universality of our experimental results.
Model Architecture Rationale.
In Appendix D.1, we discuss how does LLaVA-1.5 represents the majority of advanced MLLMs, supported by their architecture alignment. Besides, we show more empirical results on other MLLMs, Qwen-VL and LLaVA-NeXT. In Appendix D.2, we explain why DINO and PaddleOCRv2 can represent other detection models, thanks to the proposed special format. In Appendix D.3, we conduct experiments based on YOLOv5N and YOLOv11L, and inves- tigate the impact of detector accuracy. In Appendix D.4, we remove the detection data and repeat the FTBI experiment. In Appendix D.5, we unfreeze the visual encoder and repeat the experiments. In Appendix D.6, we explore the impact of a broader object detection scope.